### v3

# Leveraging Large Language Models (LLMs) to Empower Training-Free Dataset Condensation for Content-Based Recommendation

Abstract
Modern techniques in Content-based Recommendation (CBR) leverage item content information to provide personalized services to users, but suffer from resource-intensive training on large datasets. To address this issue, we explore the dataset condensation for textual CBR in this paper. The goal of dataset condensation is to synthesize a small yet informative dataset, upon which models can achieve performance comparable to those trained on large datasets. While existing condensation approaches are tailored to classification tasks for continuous data like images or embeddings, direct application of them to CBR has limitations: (1) those methods fail to synthesize discrete texts and (2) they cannot handle the interactions involving both users and items while concurrently preserving users’ preference information during the condensation process.
To bridge this gap, we investigate efficient dataset condensation for content-based recommendation. Inspired by the remarkable abilities of large language models (LLMs) in text comprehension and generation, we leverage LLMs to empower the generation of textual content during condensation. To handle the interaction data involving both users and items, we devise a dual-level condensation method: content-level and user-level. At content-level, we utilize LLMs to condense all contents of an item into a new informative title. At user-level, we design a clustering-based synthesis module, where we first utilize LLMs to extract user interests. Then, the user interests and user embeddings are incorporated to condense users and generate interactions for condensed users. Notably, the condensation paradigm of this method is forward and free from iterative optimization on the synthesized dataset, which is distinguished from previous approaches following the paradigm of nested bi-level optimization. Extensive empirical findings from our study, conducted on three authentic datasets, substantiate the efficacy of the proposed method. Particularly, we are able to approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on dataset MIND).
Index Terms:
Dataset condensation, Deep Learning, Recommendation
§§footnotetext: Equal contribution
I Introduction
Figure 1: Comparison on the paradigm of dataset synthesis between our TF-DCon and previously proposed methods in other domains.
Content-based recommenders [1, 2, 3] have made strides in mitigating the information overload dilemma, delivering items with relevant contents (i.e., news, articles, or movies) to users. Existing advanced CBR models [4, 5, 6] are trained on large-scale datasets encompassing millions of users and items. Handling these datasets significantly strains computational resources during model training. Furthermore, the recurrent retraining of recommendation models, especially for periodic updates in real-world applications, exponentially elevates costs to unsustainable levels.
The intensifying exploration of dataset condensation, also called dataset distillation, sheld light in addressing aforementioned issues. The goal of dataset condensation is to synthesize a small yet informative dataset trained on which the model can achieve comparable performance to that of a model trained on the original dataset [7, 8, 9, 10, 11, 12]. In particular, one of the most representative methods, DC [11] formulates the condensation on image data as matching gradients of network parameters between synthetic and original data. Doscond [10] continualizes the graph structure and optimizes the synthetic graph data via a one-step gradient matching mechanism. TDD [13] proposes to distill the labels of samples in a soft manner, which could be widely applied in classification tasks and conducts condensation on text embeddings to handle the discrete textual data. In spite of the promising success, dataset condensation on CBR is still unexplored.
To bridge this gap, we investigate how to effectively condense the dataset for a textual content-based recommendation. In the context of CBR, the existing condensation approaches devised in other domains [11, 8, 10, 12] are limited by following challenges:
  • How to generate discrete textual data? Current condensation methods are designed for continuous data (i.e., images and text embeddings), following a formulation of nested bi-level optimization. Under this formulation, the data is synthesized via an outer gradient, which cannot be utilized to generate discrete textual data. Therefore, it is vital for us to find a solution to generate the discrete text.
  • How to handle the interactions between users and items? In the tasks of recommendation, the goal is to utilize the information of users and items as well as the interactions between them to infer the preferences of users. However, previously proposed methods mostly involve classification tasks, which fail to handle the interaction data and condense the preference information within it [14, 15, 16].
While a recent work [14] aims at condensing collaborative filtering dataset, it is only suitable for NTK-based (Neural Tangent Kernel) architecture and fails to process textual data. Besides, CGM [16] follows the nested bi-level formulation to condense categorical data for CTR predictions and DConRec [15] designs an optimization-based method for ID-based recommendation, both of which, however, cannot generate texts as well.
Large Language Models (LLMs) show strong capacity to distill textual information [17, 18, 19], and its expertise across various fields such as medicine [20, 21], recommendation [22, 23, 24], and graph [25]. The possession of versatile nature and extensive world knowledge enables LLMs emergent abilities, such as strong text comprehension and language generation capability [17, 18]. Given its strong text comprehension and language generation capability [17, 18], we are the first to propose utilizing LLMs for processing textual content in condensation for CBR, to address the aforementioned challenges.
Based on previous discussion, we propose a LLMs-empowered Training-Free Dataset Condensation method for content-based recommendation, abbreviated as TF-DCon. Specifically, to handle the interactions and preserve the preference information in condensed dataset [14], TF-DCon is devised as a two-level solution: content-level and user-level. At content-level, we curate prompts to instruct LLMs to condense all information and contents of an item into an informative title. In user-level, we propose a clustering-based synthesis module to simultaneously generate fake users and their corresponding historical interactions, based on user interests extracted by LLMs and user embeddings. Consequently, there are two advantages of our proposed method: (1) TF-DCon omits the conventional paradigm of formulating condensation as a bi-level optimization problem [11, 7, 8]. Instead, we formulate condensation as a forward process, which is free from iteratively training datasets (as shown in Figure 1); (2) the condensed datasets generated by our method are still textual data instead of a set of text embeddings, which makes it more generalizable and flexible to train different architectures of models for CBR. Our contributions could be summarized as follows:
  • To the best of our knowledge, we are the first to investigate the problem of condensing datasets for textual CBR, where we devise a forward condensing paradigm which is different from the widely adopted bi-level optimization formulation.
  • We first propose a LLMs-empowered dataset condensation method. Specifically, we curate prompts to instruct LLMs to condense the contents of items and devise a clustering-based synthesis module to generate fake users and their historical interactions.
  • Extensive empirical findings from our study, conducted on three authentic datasets, substantiate the efficacy of the proposed method. Particularly, we are able to approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on dataset MIND). The model training on the condensed datasets is significantly faster (i.e., 5×5$\times$5 × speedup).
The rest of the paper is structured as follows. Section V reviews the related works for content-based recommendation and dataset condensation. Section II provides background for relevant preliminaries. Section III elaborates on the details of the proposed condensation method for CBR. Section IV presents the experimental results and analysis. The conclusion is obtained in section VI.
II Preliminaries
Figure 2: The proposed method in a nutshell. The LLMs-empowered content-level condensation and user interest extraction are instanced in Figure 3.
II-A Content-based Recommendation
Before elaborating on our proposed method, let us introduce some basic knowledge and a formal definition of content-based recommendation. CBR aims to infer the preference of users based on historical interactions and then recommend the contents with potential interest. In the context of CBR, the dataset $\mathcal{D}$ mainly consists of item set $\mathcal{N}$, user set $\mathcal{U}$, and click history set $\mathcal{H}$. Each item $n\in\mathcal{N}$ contains various contents, such as title, abstract, and category. Each user $u\in\mathcal{U}$ is associated with a clicking history of items $h^{(u)}\subset\mathcal{N}$. We denote the set of history $h^{(u)}$ by $\mathcal{H}$. Let $\mathcal{C}$ denote the set of clicks, where each click $c\in\mathcal{C}$ is defined as a tuple $(u,n,y)$ indicating whether user $u$ clicked on item $n$ with label $y\in\{0,1\}$. The task of CBR is to infer the preference of users on given candidate items based on their historical interactions.
II-B Dataset Condensation
The aim of dataset condensation is to synthesize a small yet informative dataset $\mathcal{S}$, trained on which the recommendation model can achieve comparable performance to the model trained on full dataset $\mathcal{D}$. Here, the synthesized dataset $\mathcal{S}$ also consists of its corresponding item set $\mathcal{N^{S}}$, user set $\mathcal{U^{S}}$ and click history set $\mathcal{H^{S}}$. Formally, the objective of the condensation could be formulated as follows:
$\min_{\mathcal{S}}\mathcal{L}(f_{\boldsymbol{\theta}^{\mathcal{S}}}(\mathcal{D})),\text{ s.t. }\boldsymbol{\theta}^{\mathcal{S}}=\underset{\boldsymbol{\theta}}{\arg\min\;}\mathcal{L}(f_{\boldsymbol{\theta}}(\mathcal{S})),$ (1)
where $\boldsymbol{\theta}$ is the parameter of recommendation model $f$ and $\mathcal{L}$ is the loss function. As we can see in Equation 1, the condensation is a bi-level problem, where the outer optimization is to synthesize the condensed dataset and the inner optimization is to train the model on dataset $\mathcal{S}$.
III Method
III-A Overview of the Proposed Method
In this section, we detail the proposed method. Due to the inherent discreteness of textual data, the conventional gradient-based dataset condensation approaches, which update the synthetic data via the outer gradient backpropagation in the bi-level formulation, is invalid for condensing textual contents. Fortunately, the emergent abilities of LLMs [17, 18], such as strong text comprehension and language generation capabilities, shed light on this issue. Hence, we utilize LLM to assist the condensation. To correspondingly address the challenges, we discuss the approaches in section III-B and section III-C. In section I, we will discuss comparison on the time complexity of different condensation methods.
As shown in Figure 2, the proposed method condenses the content-based recommendation datasets from two levels: content-level and user-level. In content-level condensation, we aim to alleviate the textual data load for each item by condensing its content (Section III-B). In user-level condensation, we aim to minimize historical redundancies by pinpointing users’ primary interests and merging users with similar preferences (Section III-C). Both content compression and user synthesis leverage the capabilities of advanced LLMs, which we’ll delve into in the subsequent sections.
III-B Content-Level Condensation
In the content-level condensation, we aim to condense all the information of each item into a succinct yet informative title. Recent studies [26, 27] reveal the exceptional power of large language models in processing textual contents. Therefore, we propose to utilize LLM to condense the contents. Specifically, we design the guiding prompt in the following format:
Hints on the format of input:
          [title]{title}, [abs]{abs},…
Instructions on the format of output:
          [new_title]{new_title}
After facilitating LLM’ comprehension on the input, the information of items will be fed into LLM to generate the condensed title. Formally, the process can be described as follows:
$n_{s}=LLM(n),$ (2)
where $n\in\mathcal{N}$ is the original contents of item, consisting of $[title]$, $[abstract]$ and $[category]$, and $n_{s}\in\mathcal{N}^{\mathcal{S}}$ is the condensed title of the item, which only contain the $[title]$. An instance of the content-level condensation is depicted in Figure 3.
III-C User-Level Condensation
In user-level condensation, we aim to condense the number of users, reducing the total size of the interactions. First, we introduce the interest extraction for each user and user encoder. Then, we propose the clustering-based synthesis module to synthesize fake users and their corresponding historical sequences. In terms of user synthesis, we apply the clustering method to the user embeddings and each cluster corresponds to a synthetic user. To synthesize interactions for each user, we merge the interactions of users, whose embedding distance and interest distance to their cluster centroids rank top-$m$ while in descending order.
III-C1 Interests Extraction and User Encoder
Interests Extraction: Each user $u$ is associated with a click history $h^{(u)}$, based on which we will utilize LLM to extract the set of interests $\mathcal{I}^{(u)}$ for this user. Specifically, we design the guiding prompt in the following format:
Hints on the format of input:
      (1){title}, (2){title}, (3){title}, …
Instructions on the format of output:
      [interests] -interest1, -interest2, …
Then, the click history $h^{(u)}$ will be fed into LLM to generate the interests of user $u$, which could be formulated as follows:
$\mathcal{I}^{(u)}=LLM(h^{(u)}),$ (3)
where $\mathcal{I}^{(u)}$ is the set of interests for user $u$.
User Encoder: For each user $u$, given the associated click history $h^{(u)}$, we first we obtain the user’s embedding as follows:
$z_{u}=f_{\theta}(h^{(u)},\mathcal{N}),$ (4)
where $\mathcal{N}$ is the set of items and $f_{\theta}$ is the well-trained user encoder [5, 2, 3].
III-C2 Clustering-Based Synthesis
User Synthesis: Given the users’ embeddings, we apply $K$-means algorithm over all of them to obtain their cluster centroids $\{c_{i}\}_{i=1}^{K_{s}}$, where $K_{s}$ is the number of synthesized users in the condensed dataset.
Historical Sequence Synthesis: Given that each cluster corresponds to a fake user $u_{s}$, we need to synthesize the corresponding historical interactions. We devise a scoring mechanism based on the distance between user interests and user embeddings to select the historical interactions for fake users.
First, we calculate the distance between the embedding of each real user $u$ and its corresponding prototype $c_{i}$ as follows:
$d^{emb}_{u}=Dis(z_{u},c_{i}),$ (5)
where $Dis(\cdot,\cdot)$ is the distance function.
Given a set of interests $\mathcal{I}^{(u)}$ for each user $u$, we encode those interests by pretrained language model (PLM) as follows:
$e_{u}=f_{pool}\left(f_{PLM}(\mathcal{I}^{(u)})\right),$ (6)
where $f_{pool}$ is the pooling function and $f_{PLM}(\cdot)$ is the pretrained language model. Based on the clusters of user embeddings, we calculate the corresponding cluster centroids of user interests $\{c^{\prime}_{i}\}_{i=1}^{K_{s}}$ by averaging the user interests in each cluster. Given the interest embeddings and their centroids, we can calculate the distance between them as follows:
$d^{int}_{u}=Dis(e_{u},c^{\prime}_{i}).$ (7)
Combining Equation 5 and Equation 7, we have:
$d_{u}=d^{emb}_{u}+\alpha\cdot d^{int}_{u},$ (8)
where $\alpha$ is a hyperparameter and $d_{u}$ is defined as the selection score. Within each cluster, we employ an ascending ordering of users based on their selection scores. We subsequently curate the historical interactions for a synthetic user by merging the interaction histories of the top-$m$ users, which could be formalized as follows:
$h_{s}^{(u_{s})}=\cup_{i=1}^{m}h^{(u_{i})},$ (9)
where $d_{u_{i}}$ ranks top-$m$ within its corresponding cluster and $h^{(u)}\in\mathcal{H}$ is the historical interactions of $u$ in the original dataset.
Figure 3: An instance of content-level condensation and user interest extraction.
III-D Discussion
TABLE I: Comparison of Computational Complexity.
Category Methods Inner Outer Total
Kernel-based Distill-CF [14] 1 $\tau_{o}$ $\tau_{o}$
Matching-Based DC [11] $\tau_{i}$ $\tau_{o}$ $\tau_{i}\cdot\tau_{o}$
Doscond [10] 1 $\tau_{o}$ $\tau_{o}$
Training-Free TF-DCon (Ours) 1 1 1
Complexity Comparison. The condensation methods are categorized into two types [28]: matching-based and kernel-based. The first type generates synthetic datasets via gradient matchings between the model trained on small-condensed dataset and those trained on the large-real dataset [11, 10], which typically involve a complex bi-level optimization. In the inner optimization, the models are trained while the datasets are optimized via outer gradients. In contrast, the kernel-based methods reduce the problem into a single-level optimization, which is always free from model training during condensation. If we denote the number of outer-iteration as $\tau_{o}$ and the number of inner-iteration as $\tau_{i}$, we can respectively obtain the computational complexity for each and we compare our proposed method with several representative condensation approaches. As shown in Table I, we can find that our proposed method is the most efficient since it follows a forward paradigm, thus free from iterative optimization on the synthetic dataset.
IV Experiments
In this section, we conduct experiments to evaluate TF-DCon, mainly aiming to answer the following research questions: RQ1: how well the performance can be achieved when trained on datasets condensed by TF-Con? RQ2: how generalizable is the datasets condensed by TF-Con across different recommendation models?
IV-A Experimental Settings
IV-A1 Datasets
To evaluate the performance of our method, we condense the training set of three real-world datasets: MIND [1], Goodreads [29], and MovieLens [30]. The statistics of these datasets are shown in Table II. We split the datasets into $80\%/10\%/10\%$ for training/validation/test. The condensation is mainly conducted on the training sets of datasets and during the testing phase, we still use the users and contents of items from the original datasets.
TABLE II: Dataset statistics. “avg. tok.” denotes the average number of tokens in the content of each item and “avg. his.” denotes the average length of user histories.
Datasets MIND GoodReads MovieLens
# Items 65,238 16,833 1,682
avg. tok. 45.42 35.37 34.80
# Users 94,057 23,089 943
avg. his. 14.98 7.81 4.94
# pos 347,727 273,888 48,884
# neg 8,236,715 485,233 17,431
Density 0.0057% 0.0705% 3.0820%
IV-A2 Baselines
Due to the limited studies on condensation for CBR, we compare our proposed methods with two baselines implemented by ourselves: (i) Random. In the setting of overall performance comparison (Table III), we randomly select a certain portion of users and randomly select a certain portion of tokens to be the contents for each item. In the setting of user-level condensation (Table VII), it is implemented via randomly selecting certain ratios of users while the contents of items remain the same as those in original datasets. In the setting of content-level condensation (Table VI), we implement it by randomly selecting tokens for each item. (ii) Majority. In the overall performance comparison, we implement this method by sampling the users with a large number of interactions and compressing the contents of items by randomly sampling the tokens.
IV-A3 Evaluation Protocol
To evaluate the quality of our method, we test the performance of recommendation models trained on the condensed datasets. In this evaluation, we employ three prevalent content-based recommendation models, i.e., NAML [2], NRMS [5] and Fastformer [3]. They are all designed in a hierarchical manner, i.e., using a content encoder to capture item representations in the history sequence, and a user encoder to fuse the historical item representations into a unified user representation. To be specific, the evaluation is threefold: (1) condense the datasets, (2) train the recommendation models on the condensed datasets, and (3) test the performance of the models. For the evaluation metrics, we follow the common practice [5, 2] to adopt the widely used metrics, i.e., NDCG@$K$ and Recall@$K$, abbreviated as N@$K$ and R@$K$ in the tables, respectively. In this work, we set $K=1,5$ for evaluation on Goodreads and MovieLens, and set $K=5,10$ for MIND. To compare the performance of models trained on condensed datasets and original datasets, we average the percentages of metrics relative to the metrics trained on original datasets. We denote the percentages by “Quality”.
IV-A4 Implementation Details
During training, we employ Adam [31] operator with a learning rate of 5e-3 for three datasets. We select title, abstract, and category feature of each piece of news as the item content for the MIND dataset, and select book or movie name and description feature of each book or movie as the item content for the Goodreads and MovieLens dataset. We concatenate multiple item contents into a single sequence before feeding them to the recommenders. For all three backbone models, we set the embedding dimension of content encoders to 256, the embedding dimension of user encoders to 64, and the negative sampling ratio to 4. We tune the hyperparameters of all base models to attain optimal performance. We average the results of five independent runs for each model. All experiments are conducted on a single NVIDIA GeForce RTX 3090 device.
IV-B Overall Comparison
TABLE III: Overall Performance Comparison on Condensed Datasets. “Quality” denotes the achieved ratio of performance when compared to those trained on original datasets. The detailed condensation ratio could be found in Table IV.
Datasets MIND1, $r$=5Goodreads, $r$=27MovieLens, $r$=26%
Rec Model Metrics Random Majority TF-DCon Original Random Majority TF-DCon Original Random Majority TF-DCon Original
N@1 0.2871 0.2854 0.3071 0.3176 0.5197 0.5057 0.5411 0.6462 0.8241 0.8397 0.8484 0.828
N@5 0.3470 0.3466 0.3691 0.3783 0.7943 0.7884 0.8033 0.8475 0.8251 0.8367 0.8494 0.831
R@1 0.4016 0.4002 0.4377 0.4534 0.4520 0.4326 0.4704 0.5635 0.1785 0.1886 0.1873 0.1752
R@5 0.5670 0.5697 0.6150 0.6270 0.9983 0.9986 0.9984 0.9989 0.7274 0.7323 0.7475 0.7399
NAML Quality 90.28% 90.15% 97.22% 100.00% 88.57% 87.01% 90.49% 100.00% 99.75% 102.18% 103.15% 100.00%
N@1 0.2625 0.2631 0.2997 0.3009 0.5399 0.5094 0.5453 0.6439 0.8105 0.8294 0.8149 0.8178
N@5 0.3225 0.3225 0.3597 0.3608 0.8017 0.7901 0.8054 0.8476 0.8227 0.8334 0.8371 0.8253
R@1 0.3750 0.3793 0.4279 0.4325 0.4704 0.4344 0.4751 0.5629 0.1729 0.1821 0.1798 0.1757
R@5 0.5414 0.5445 0.6000 0.6042 0.9982 0.9990 0.9985 0.9993 0.7325 0.7339 0.7446 0.7321
NRMS Quality 88.23% 88.66% 99.38% 100.00% 90.47% 87.37% 91.01% 100.00% 99.31% 101.57% 101.28% 100.00%
N@1 0.2815 0.2736 0.3022 0.3057 0.5420 0.5165 0.5548 0.6556 0.7886 0.8105 0.8251 0.7915
N@5 0.3425 0.3350 0.3637 0.3645 0.8028 0.7934 0.8092 0.8529 0.8254 0.8318 0.8429 0.8145
R@1 0.3944 0.3804 0.4334 0.4365 0.4725 0.4414 0.4851 0.5745 0.1738 0.1799 0.1836 0.166
R@5 0.5631 0.5515 0.6096 0.6144 0.9983 0.9991 0.9986 0.9990 0.739 0.7373 0.7465 0.7279
Fastformer Quality 92.01% 89.58% 99.29% 100.00% 89.74% 87.16% 90.97% 100.00% 101.80% 103.55% 105.22% 100.00%
  • 1
    For the dataset MIND, we report $N@5,N@10$ and $R@5,R@10$, which is also the same in the subsequent tables of ablation studies.
TABLE IV: Comparison between condensed/sampled datasets and original datasets. We use “ORI.”, “RD.” and “MJ.” to represent the statistics of original, random sampled, and majority-selected dataset. “Item” and “User” denote the condensation information of item contents and users, respectively.
Datasets MIND Goodreads MovieLens
Methods OR. RD. MJ. Ours OR. RD. MJ. Ours OR. RD. MJ. Ours
Item avg. tok. 45.42 16.73 16.73 16.73 35.37 16.01 16.73 16.01 34.80 15.26 16.73 15.26
Size (KB) 10,384 4,095 4,095 4,095 2,189 993 993 993 232 121 121 121
Ratio 100% 39% 39% 39% 100% 45% 45% 45% 100% 52% 52% 52%
User # users 94,057 9,405 9,405 9,405 23,089 4,617 2,350 4,617 943 47 47 47
Size (KB) 106,614 2,124 5,733 2,139 8,704 1,791 1,977 1,957 532 68 232 81
Ratio 100% 2% 5% 2% 100% 21% 23% 22% 100% 13% 44% 15%
Overall Size (KB) 116,998 6,219 9,828 6,234 10,893 2,784 2,970 2,950 764 189 353 202
Ratio 100% 5% 8% 5% 100% 26% 27% 27% 100% 25% 46% 26%
To validate the effectiveness of TF-DCon (RQ1), we measure the recommendation performance of CBR models trained in the condensed datasets and their training efficiency. Specifically, we train the recommendation models (i.e., NAML, NRMS, Fastformer) on the condensed datasets and evaluate their performance on the test set of the original datasets. The results are reported in Table III and the detailed condensation ratios are shown in Table IV. Further, we also evaluate the training efficiency on the condensed datasets and the results are summarized in Figure 4.
IV-B1 Overall Performance Comparison
The proposed TF-DCon achieves better performance than the two baselines on three datasets. Notably, as shown in Table III, we approximate over $97\%$ of the original performance with only $5\%$ data on dataset MIND across three different recommender models. From Table IV, we observe that the amazing condensation ratio comes from two parts: reducing the size of contents by $61\%$ (item side) and reducing the size of historical interactions and users (user side) by $98\%$. We observe similar performance on datasets Goodreads and MovieLens. Interestingly, when trained on condensed MovieLens, we may achieve a slightly better performance than those trained on the original dataset. We attribute this to the small size of the original MovieLens dataset since the condensed datasets may serve as a refined and denoised alternative to the original one.
From Table IV, we can find that in each dataset, the size of the user part (i.e., users’ historical interactions) is much larger than the size of the item part (contents of items). Consequently, the predominant size reduction stems from the condensation of the user part. We attribute this promising reduction to the common interests among different users, thus rendering that method of generating representative fake users to replace a variety users reasonably and effective. The condensation ratio of user-side on Goodreads and MovieLens dataset is less promising than the ratio on MIND. This discrepancy is ascribed to the fact that smaller datasets inherently exhibit fewer redundant interests among distinct users, consequently leading to a less pronounced reduction ratio.
IV-B2 Training Efficiency
[…]
005555101010101515151520202020252525253030303035353535404040400.260.260.260.260.270.270.270.270.280.280.280.280.290.290.290.290.30.30.30.30.310.310.310.310.320.320.320.32Training Time (min)NDCG@5
(a) MIND
00202020204040404060606060808080801001001001000.740.740.740.740.760.760.760.760.780.780.780.780.80.80.80.80.820.820.820.820.840.840.840.840.860.860.860.86Training Time (min)NDCG@5
(b) Goodreads
Figure 4: Training efficiency on original and condensed datasets.
Here, we compare the training time and performance of NAML, NRMS and Fastformer when trained on the original datasets and condensed datasets by TF-DCon . As shown in Figure 4, the solid lines represent the training curves on condensed datasets and the dotted line represents original datasets, where we can observe that the training on condensed datasets converges much faster than the training on original datasets (i.e., up to 5×5$\times$5 × speedup). Typically, the model convergences on condensed datasets are reached with around $6$ minutes while the training on original datasets needs around or more than $30$ minutes. For the performance, we find that those trained on condensed datasets can exhibit comparability with regard to the metric $NDCG@5$.
IV-C Generalization Across Recommender Models
IV-C1 Generalizability of Condensed Datasets
To answer RQ2, we test the performance of CBR models with different architectures trained on the datasets condensed by one CBR model. Specifically, we employ NAML [2] as the backbone model for condensation and test the performance of NAML [2], NRMS [5] and Fastformer [3] trained on the condensed datasets. The results are reported in Table III. Furthermore, we also test the generalizability of content-level and user-level condensation in the ablation study for TF-DCon , as shown in Table VI and Table VII. From the results, we can observe that the datasets condensed by TF-DCon exhibit good generalization on different CBR models. We ascribe this capability to textual data, as texts exclusively encapsulate semantic information while excluding architectural details found in text embeddings.
IV-C2 Versatility of TF-DCon
The proposed TF-DCon is highly flexible since we can adopt different CBR models as the user encoder in the condensation method. We investigate the performances of different models on datasets condensed by TF-DCon with different recommendation models. We utilize NAML, NRMS, and Fastformer as the user encoder for both condensation and evaluation, respectively. We choose dataset MIND to report results in Table V, where we can observe the strong generalization ability of condensed datasets on other models. We can also observe that NAML can outperform the other two models when trained on the most condensed dataset as well as the original dataset.
TABLE V: TF-DCon with different recommender models for condensation. The datasets condensed by different recommender models exhibit similar training performance across different recommender models, verifying the versatility of our method.
Train Set TF-DCon ${}_{NAML}$ TF-DCon ${}_{NRMS}$ TF-DCon ${}_{Fast.}$ Original
Metrics N@5 R@5 N@5 R@5 N@5 R@5 N@5 R@5
NAML 0.3071 0.4377 0.3066 0.4396 0.3060 0.4368 0.3176 0.4534
NRMS 0.2997 0.4279 0.2905 0.4256 0.2924 0.4184 0.3009 0.4325
Fastformer 0.3022 0.4334 0.3094 0.4409 0.3056 0.4343 0.3057 0.4365
IV-D Ablation Study
To examine how user-level condensation and content-level condensation affect the overall performance, we performed an ablation study on them, respectively. We examine the performance of models trained on datasets generated by only user-level condensation or content-level condensation. Specifically, content-level condensation only condense the contents of items and the historical interactions. The results are reported in Table VI, where we observe that only condensing the item contents can also achieve comparable performance. The condensation ratio is relatively not promising and we attribute this to the information capacity of the contents which restricts the size of textual data.
For user-level condensation, we only condense the users into fake users and their corresponding interactions while the contents are preserved. As shown in Table VII, we can find that the condensation ratio is astonishingly good while we can still achieve good performance. Especially, on dataset MIND, we can achieve over 97% performance of the original datasets with 98% size reduction of the user-side dataset (i.e., the historical interactions of users ) when tested on Fastformer. When tested on NAML and NRMS, the performance can also exceed over 93% of those trained on the original dataset. Besides, the condensation ratios on Goodreads and MovieLens are also good with a comparable performance. We attribute the promising condensation ratios to similar interests among different users. Therefore, when we synthesize a fake user to represent a group of users with common interests, we can provide adequate information for model training.
[…]
Figure 5: Varying the hyperparameter $\alpha$ in Eq. 8.
TABLE VI: Ablation Study on Content-Level Condensation.
Datasets MIND, $r=39\%$ Goodreads, $r=45\%$ MovieLens, $r=52\%$
Rec Model Metrics Random TF-DCon Original Random TF-DCon Original Random TF-DCon Original
NAML N@1 0.3059 0.3126 0.3176 0.5807 0.6359 0.6462 0.7959 0.8382 0.8280
N@5 0.3678 0.3725 0.3783 0.8204 0.8434 0.8475 0.8208 0.8367 0.8310
R@1 0.4396 0.4457 0.4534 0.5053 0.5543 0.5635 0.1665 0.1811 0.1752
R@5 0.6167 0.6169 0.6270 0.9988 0.9989 0.9989 0.7341 0.7374 0.7399
Quality 97.39% 98.39% 100% 95.06% 99.23% 100% 97.79% 100.75% 100%
NRMS N@1 0.2902 0.3050 0.3009 0.6320 0.6668 0.6439 0.8017 0.8207 0.8178
N@5 0.3525 0.3626 0.3608 0.8421 0.8561 0.8476 0.8250 0.8285 0.8253
R@1 0.4205 0.4387 0.4325 0.5521 0.5832 0.5629 0.1724 0.1759 0.1757
R@5 0.5998 0.6041 0.6042 0.9991 0.9990 0.9993 0.7329 0.7354 0.7321
Quality 97.23% 99.30% 100% 97.43% 98.34% 100% 99.26% 100.38% 100%
Fastformer N@1 0.2988 0.3025 0.3057 0.6296 0.6499 0.6556 0.7828 0.7974 0.7915
N@5 0.3599 0.3643 0.3645 0.8408 0.8506 0.8529 0.8086 0.8159 0.8145
R@1 0.4321 0.4359 0.4365 0.5495 0.5701 0.5745 0.1653 0.1699 0.1660
R@5 0.6067 0.6129 0.6144 0.9991 0.9991 0.9990 0.7223 0.7266 0.7279
Quality 98.63% 99.67% 100% 97.96% 99.60% 100% 99.16% 100.40% 100%
TABLE VII: Ablation Study on User-Level Condensation.
Datasets MIND, $r=2\%$ Goodreads, $r=22\%$ MovieLens, $r=15\%$
Rec Model Metrics Random TF-DCon Original Random TF-DCon Original Random TF-DCon Original
NAML N@1 0.2886 0.2964 0.3176 0.5259 0.5391 0.6462 0.8154 0.8338 0.8280
N@5 0.3474 0.3528 0.3783 0.7970 0.8034 0.8475 0.8350 0.8470 0.8310
R@1 0.3995 0.4273 0.4534 0.4570 0.4736 0.5635 0.1776 0.1879 0.1752
R@5 0.5627 0.5887 0.6270 0.9983 0.9983 0.9989 0.7457 0.7468 0.7399
Quality 89.97% 93.75% 100% 90.91% 92.09% 100% 99.98% 101.61% 100%
NRMS N@1 0.2702 0.2834 0.3009 0.5488 0.5559 0.6439 0.8132 0.8207 0.8178
N@5 0.3281 0.3449 0.3608 0.8066 0.8099 0.8476 0.8208 0.8338 0.8253
R@1 0.3796 0.4159 0.4325 0.4793 0.4851 0.5629 0.1738 0.1802 0.1757
R@5 0.5413 0.5927 0.6042 0.9984 0.9984 0.9993 0.7291 0.7373 0.7321
Quality 89.45% 96.38% 100% 92.78% 93.31% 100% 99.45% 100.83% 100%
Fastformer N@1 0.2880 0.2978 0.3057 0.5564 0.5614 0.6556 0.8047 0.8251 0.7915
N@5 0.3469 0.3568 0.3645 0.8094 0.8113 0.8529 0.8260 0.8412 0.8145
R@1 0.3931 0.4255 0.4365 0.4860 0.4902 0.5745 0.1740 0.1815 0.1660
R@5 0.5575 0.5952 0.6144 0.9985 0.9983 0.9990 0.7359 0.7460 0.7279
Quality 92.12% 97.34% 100% 92.48% 92.84% 100% 101.63% 103.76% 100%
IV-E Further Investigation
In this subsection, we further investigate how different values of hyperparameter $\alpha$ affect the quality of the condensed datasets. The hyperparameter $\alpha$ is utilized to balance the influence of user interests and user embeddings on the user synthesis. The larger value of $\alpha$ denotes larger influence of user interests on the synthesis of fake users. As shown in Figure 5, we can find that the best choice for $\alpha$ with three different recommendation models is smaller than 1 but larger than 0, which verifies that the effectiveness of selecting/merging users with similar interests for the synthesis.
V Related Work
V-A Content-Based Recommendation
Content-based recommendation (CBR) involves a wide range of information in different formats: text, image, video, social networks, etc [32]. In this paper, we mainly focus on textual contents, i.e., news, books, and movie reviews.
In recent years, textual CBR has garnered substantial attention and has been the subject of comprehensive investigation within both industrial and academic domains. Models based on neural networks [4, 5] have been proposed to learn the representations of users and items via CNN [33] or attention mechanism [34]. Thereafter, the advance of pretrained language models like BERT [35] exhibited promising performance on textual content recommendation [36, 37]. As those methodologies depend on training with extensive-scale datasets, their exceptional performance comes at a high computational cost. The recent emergence of inquiries into dataset condensation and dataset distillation presents a promising avenue for tackling this concern [7, 11]. The inherent concept underlying dataset condensation and dataset distillation revolves around the synthesis of a compact dataset for model training, with the objective of enabling models trained on this condensed dataset to achieve performance levels comparable to those trained on the complete dataset. Nevertheless, no extant research has delved into the realm of condensation specifically within the context of CBR. To bridge this critical gap, we introduce an innovative and efficacious framework tailored to the task of dataset condensation for CBR.
V-B Dataset Condensation
It is notorious that training neural networks on large datasets is prohibitively expensive. To mitigate this issue, plenty of works on dataset distillation/dataset condensation have been proposed to synthesize small datasets to replace the large ones, such as GCond [8], DC [11] and SFGC [9]. Those methods always formulate the condensation process as a bi-level optimization problem. The condensation methods have been explored on images [11], graph [8] and text embeddings [12]. Various methods have been proposed to improve the quality of condensation and optimization efficiency. For instance, DC [11] proposed gradient matching, achieving a comparable performance under a high compression ratio, and DosCond [10] proposed one-step update strategy, which significantly improves the optimization efficiency. Condensation on textual data is conducted based on the text embeddings or soft labels [38, 39, 40, 13, 12]. Further exploration extends condensation methods to image-text retrival [41]. Recently, DConRec [15] is proposed to condense the datasets for ID-based recommendation and CGM [16] is proposed to handle the datasets for CTR prediction with side information. Distill-CF [14] proposes a condensation method tailored for NTK-based models.
However, rare attention has been paid to the domain of CBR. To address this disparity, we posit the consolidation of the CBR dataset through the utilization of Large Language Models (LLM). Unlike earlier methodologies, our dataset condensation approach is devoid of training requirements, leading to a substantial reduction in condensation expenses. Furthermore, our condensation procedure centers on textual content, thereby diverging from those text condensation methods that hinge on text embeddings.
VI Conclusion
Advanced content-based recommendation models rely on training with large datasets, which is highly expensive. To solve this problem, in this paper, we investigate how to condense the dataset of content-based recommendation into a small synthetic dataset, trained on which models can achieve comparable performance with those trained on the full dataset. Consequently, we propose a novel method TF-DCon that utilizes LLMs to condense contents of items and users, where users and their corresponding historical interactions are condensed via a curated clustering-based synthesis module. To the best of our knowledge, this is the first work on condensing datasets for content-based recommendation and the first condensation method that is free from iterative optimizing synthetic data. The experimental results on multiple real-world datasets verify the effectiveness of our proposed methods. Notably, we approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on dataset MIND). The model training on the condensed datasets is significantly faster (i.e., 5×5$\times$5 × speedup).